OpenStudy (anonymous):

A spherical model of planet earth has a radius of 2 ft. What is the approximate volume of the model of planet earth? Use 3.14 to approximate pi and round your answer to two decimal places is the answer 33.49 or 33.51 @inowalst @Destinymasha

OpenStudy (inowalst):

\[V=l \times w \times h\]

OpenStudy (igreen):

No..that's for a rectangular prism..

OpenStudy (igreen):

\(V = \dfrac{4}{3} \pi r^3\) \(V = \dfrac{4}{3} (3.14) (2)^3\) Multiply 4/3 * 3.14 * 2 * 2 * 2

OpenStudy (anonymous):

which one is it @iGreen

OpenStudy (anonymous):

I did that and I get 33.49 is that correct @iGreen

OpenStudy (igreen):

Yep, you got it.

OpenStudy (anonymous):

thnks